# Multi Task Inverse Reinforcement Learning for Common Sense Reward

### v1

@@ 6 Conclusions
In this work, we addressed the important question of how can we get rewards capable of guiding reinforcement learning agents to exhibit desired behavior within our environment. An important step in achieving this goal is to disentangle the task-independent reward, which we name common-sense reward from the task reward. We believe this simple observation can have an important impact as it allows us to focus on learning how the agent should behave and not on what the agent should be doing.
Furthermore, our experiments show that not only does our new framework allow us to easily combine expert observations from different tasks, but that learning from a variety of different tasks is a key to learning meaningful reward functions.
Finally, we observe that current IRL methods, while successful in imitation learning, do not learn a proper reward function. We advocate for further work in this direction, as the reward function can be of great importance and not just an auxiliary for imitation learning.
Appendix A Additional experimental results for Single Task Experiment
@@
In section 5.2 we explained about figure 3, showing that during the IRL process the trained agent successfully learned the desired common-sense behavior. However, a new agent training on the learned reward does not achieve the same desired behavior. Here we will show those results on two more tasks. In the paper, figure 3 shows graphs for the ”reach-wall-v2” task. Results fot ”drawer-close-v2”, and ”button-press-topdown-wall-v2” are shown in figures 7 and 8, respectively.
Figure 7: Single task cs-reward: Velocity and action norm during IRL, and during RL training with the learned cs-reward, on ”drawer-close” task.
Figure 8: Single task cs-reward: Velocity and action norm during IRL, and during RL training with the learned cs-reward, on ”button-press-topdown-wall” task.
Appendix B Additional experimental results for Multi Task Experiment
@@
In section 5.4, we showed in a scatter plot the correlation between the learned cs reward and the ground truth reward. We showed it for the velocity case, and here in figure 9 we will show it also for the action norm case. The Correlation coefficient between ground truth reward and cs reward is 0.81. In the action norm case, similarly to the velocity case, we can see that the cs reward is highly correlated with ground truth reward.
Figure 9: Scatter plot of ground-truth VS learned cs-reward for MT-CSIRL on velocity experiment. The reward is an agent trained on an unseen task.
Appendix C More Experiments’ Details
In Alg. 1, we explain the MT-CSIRL training process. The inputs of the algorithm are:
  • Expert demonstrations $\mathcal{D}_{t_{1}}...\mathcal{D}_{t_{T}}$, which in our experiments set to 10K demonstrations for each task.
  • Discriminator updates per iteration $D_{\text{updates}}$ is set to 10
  • Number of tasks $T$: in the multi target experiment is set to 5, and in the multi task experiment is set to 5 targets * 6 tasks.
  • For the RL generator, we used Soft Actor Critic, with HP as in (Yu et al., 2019)
  • SAC was updates per iteration $G_{\text{updates}}$,is set to 20.
C.1 Common Sense Reward with Sawyer Robot
In our experimental framework, we utilize the Meta-World benchmark, which incorporates the Sawyer robot, a collaborative robotic arm known for its high precision and 7-degree-of-freedom. This setup enables us to simulate and evaluate complex tasks, and to define a common sense, which we showed in our paper how we learned ans transfer it between tasks. Our velocity ground truth reward, was calculated on the ”end effector” part of the sawyer robot shown in figure 10. The end effector located at the end of the arm, there is an end effector, which can be equipped with different tools or grippers, depending on the task. In simulation environments like MetaWorld, the end effector is often a gripper used for tasks like picking and placing objects.
The location of the end effector is given us by both Mujoco and Meta-worlds observation space. In each step we have an access to the end effector of $s$ ans $s^{\prime}$, and we use those loations to calculate the velocity, as explained in 5.1
We introduce two common sense behaviors, as shown in section 5.1. In the velocity case, we calculate the velocity of the Sawyer’s end effector. Our reward function is maximized when the end effector velocity get closer to $C_{v}$. In our experiments, we set $C_{v}$ to be 0.8. As for the action norm, our ground truth reward is maximized when the $l_{1}$ of the action space, gets closer to $C_{n}$.In our experiments, we set $C_{n}$ to be 3. It is important to mention that action space is the same for all meta worlds tasks.
@@ C.1 Common Sense Reward with Sawyer Robot
Figure 10: Visualization of Sawyer Robot: The end effector which is the part we calculated the velocity on, marked with an arrow.

### v2

@@ 6 Conclusions
In this work, we addressed the important question of how to learn rewards capable of guiding reinforcement learning agents to exhibit desired behavior within our environment. An important step in achieving this goal is to disentangle the task-independent reward, which we name common-sense reward from the task reward. We believe this simple observation can have an important impact as it allows us to focus on learning how the agent should behave and not on what the agent should be doing. Furthermore, our experiments show that our framework allows us to easily combine expert observations from different tasks and that learning from a variety of different tasks is a key to learning meaningful reward functions. Finally, we observe that current IRL methods, while successful in imitation learning, still struggle to learn a proper reward function. We advocate for further work in this direction, as the reward function can be more than an auxiliary for imitation learning.
7 Limitation and Broader Impact
In this work, we experiment with simple synthetic common-sense rewards. These rewards are informative for this study, as they show how current methods do not learn a useful or transferable reward even in this basic setting. However, further research is required with a more realistic common-sense reward. This will require designing a novel and complex RL benchmark and is beyond the scope of this work. Regarding broader impact, this work impacts the way we train agents with better alignment with desired behavior.
Appendix A Additional Experiment Results
This section includes additional experiments’ results, introduced in the order presented in the Experiment section 5 of the paper.
A.1 Additional experimental results for CS-Reward Trained on a Single Task
@@
In section 5.1 we explained about figure 2, showing that during the IRL process the trained agent successfully learned the desired common-sense behavior. However, a new agent training on the learned reward does not achieve the same desired behavior. Here we will show those results on two more tasks. In the paper, figure 2 shows graphs for the "reach-wall-v2" task. Results fot "drawer-close-v2", and "button-press-topdown-wall-v2" are shown in Fig. 4 and 5 respectively.
(a) Velocity CS reward
(b) Action norm CS reward
Figure 4: Single task cs-reward: Visualization of the rewards during the IRL process (IRL Agent), an RL agent with the learned cs-reward from the IRL process (Transferred), and a baseline RL agent without any common sense component (SAC). The red horizontal line represents the target value in the ground truth reward, see Eq. 10 and 11. This experiment was conducted on the button-press-topdown-wall setup task.
(a) Velocity CS reward
(b) Action norm CS reward
Figure 5: Single task cs-reward:Visualization of the rewards during the IRL process (IRL Agent), an RL agent with the learned cs-reward from the IRL process (Transferred), and a baseline RL agent without any common sense component (SAC). The red horizontal line represents the target value in the ground truth reward, see Eq. 10 and 11. This experiment was conducted on the coffee-button setup task.
In Fig. 2 (c), we present a scatter plot of the learned cs-reward versus the ground-truth cs-reward, which is a result of the experiment in section 5.1. The correlation results in Fig. 2 are for the velocity experiment.
@@
Here we show the correlation between Ground-Truth & Learned CS-Reward Trained on a Single Task, for the action norm experiment. with the scatter plot visualization in Fig. 6. Similarly to the velocity experiment, the action norm experiment shows no correlation between the learned reward and the ground-truth reward, which further demonstrates that the IRL fails to capture the desired reward.
Figure 6: Action Norm Scatter plot of GT & CS-Reward Learned using Single Task
A.2 Additional experimental results for Learning CS-Reward with Multiple Targets
In section 5.2, we show that training our IRL method on expert demonstrations from different targets allows learning a reward that can train new agents on unseen targets, but it does not transfer to novel tasks. All agents achieve a 100% success rate, so we only show results for the cs-reward.
Here we show, in Table 3, the results of the velocity common-sense reward in the multiple target setup. We train our MT-CSIRL method with experts on different targets for the same task and test these learned rewards on new targets and tasks.
The diagonal elements show our reward transfers well to new targets for the seen task, with the baseline results for agents trained only on the task reward in parenthesis. On the off-diagonal elements indicate poor transfer to novel tasks, highlighting the limitations of our cs-rewards.
As an additional baseline for the multiple targets experiment, we report the velocity when the task is learned using only the task-specific reward, without the learned velocity cs-reward, with the results presented in parentheses.
reach-wall BP Topdown-wall drawer-close push-back coffee-button
reach-wall $\mathbf{0.93}(0.37)$ $0.40$ $0.43$ $0.40$ $0.25$
BP Topdown-wall $0.36$ $\mathbf{0.82}(0.24)$ $0.50$ $0.31$ $0.21$
drawer-close $0.34$ $0.48$ $\mathbf{0.89}(0.26)$ $0.35$ $0.24$
push-back $0.25$ $0.25$ $0.20$ $\mathbf{0.88}(0.28)$ $0.22$
coffee-button $0.28$ $0.32$ $0.35$ $0.38$ $\mathbf{0.91}(0.26)$
Table 3: Velocity cs-reward. Each row represents a different task for MT-CSIRL training, each column represents a task for RL training and evaluation. The Numbers represent the ratio between the agents’ velocity and its target value, closer to one is better. In the diagonal, Velocity reward for training solely on the task reward appears in parentheses
A.3 Additional experimental results for Learning CS-Reward with Multiple Tasks
In section 5.3, we show that training our IRL method process on experts’ trajectories from multiple tasks allows us to learn a cs-reward that can transfer to new, unseen tasks.
We perform our MT-CSIRL method again, training each expert on a different Meta-world task. In Fig. 7 we visualize a scatter plot of the Ground Truth cs-reward versus our learned reward on a new unseen task for the Action Norm case.
The scatter plot shows a high correlation with a correlation coefficient of 0.86. This indicates that in the velocity case too, our agent successfully learned the desired reward function.
Figure 7: Action Norm Scatter plot of GT & CS-Reward Learned using MT-CSIRL
A.4 Additional experimental results for Learning CS-Reward and Task-Reward with Multiple Tasks
In section 4.4, an explanation of the extension to unknown task reward method is explained, and in section 5.4, we implement the extension to this method. We assume expert demonstrations from T tasks with unknown task-specific rewards. Using MT-CSIRL+LT, we train the IRL process on multiple tasks, resulting in a learned cs-reward and task-specific rewards for each task.
In this setup, for each task $t$ we have expert demonstration, and we do not have the explicit task reward the expert’s demonstrations maximizing. To learn both cs-reward, and task-specific reward, we simultaneously learn per-task reward functions from each expert’s demonstrations and a shared common sense reward from all experts. For each expert, we train a task discriminator to learn the task-specific reward.
We use the learned cs-reward and the learned task-specific reward, and we train from scratch an agent that needs to maximize both cs-reward and task-specific reward. Results for this evaluation are shown in Table 4
Action Norm Velocity
env Action Norm Ratio Success-rate Velocity Ratio Success-rate
window-open $0.88$ $87.80\pm 1.77$ $0.90$ $88.7\pm 1.72$
reach $0.91$ $91.25\pm 1.67$ $0.89$ $84.2\pm 2.14$
plate-slide $0.90$ $90.45\pm 3.01$ $0.88$ $83.3\pm 2.32$
faucet-open $0.89$ $88.69\pm 2.42$ $0.88$ $92.3\pm 2.33$
coffee-button $0.91$ $91.81\pm 2.71$ $0.90$ $87.5\pm 1.23$
door-close $0.90$ $86.92\pm 2.60$ $0.89$ $90.0\pm 1.41$
Table 4: SAC Agent performance, with Learned Common Sense Reward and Learned Task Reward using MT-CSIRL+LT process. No transformation - learning from scratch the task whose rollouts are seen in IRL process
Algorithm 2 MT-CSIRL+LT
1:Input: Expert demonstrations $\{\mathcal{D}_{t_{1}},\ldots,\mathcal{D}_{t_{T}}\}$, number of iterations NN, discriminator updates per iteration $D_{\text{updates}}$, number of tasks $T$, generator updates per iteration $G_{\text{updates}}$
2:Initialize: Policy parameters $\omega_{t}$, common sence discriminator parameters $\theta_{cs}$, and task discriminator parameters $\phi_{t_{i}}$
3:for $i=1$ to $N$ do
4:  for $t=1$ to $T$ do
5:   for $j=1$ to $D_{\text{updates}}$ do
6:     Sample trajectories $\tau_{t}\sim\pi_{\omega_{t}}$
7:     Update CS-discriminator parameters $\theta_{cs}$ using gradient ascent
8:     Update Task-specific discriminator parameters $\phi_{t_{i}}$ using gradient ascent
9:   end for
10:   for $k=1$ to $G_{\text{updates}}$ do
11:     Sample trajectories $\tau_{t}\sim\pi_{\omega_{t}}$
12:     Update $\omega_{t}$ using $\bar{r}_{t_{i}}+\alpha\cdot r_{\text{CS}}$
13:   end for
14:  end for
15:end for
Appendix B Experimental Details
Environments We evaluate our experiments on the Meta-world benchmark. All tasks are performed by a simulated Sawyer robot. The action space consists of the 3D change of the end-effector and the normalized torque of the gripper fingers, ranging from -1 to 1. The robot either manipulates one object with a variable goal or two objects with a fixed goal. The observation space is a 39-dimensional 6-tuple of the 3D positions of the end-effector, gripper openness, positions and quaternions of two objects, and the goal. If no second object or goal, corresponding values are zeroed out.
Figure 8: Visualization of train and test tasks for our three experimental setups: Single Target, experiments in section 5.1, experiments in section Multi Target 5.2, and Multi Task, experiments in sections 5.3, 5.4
Common Sense Reward - Technical Details In our experimental framework, we utilize the Meta-World benchmark, which incorporates the Sawyer robot, a collaborative robotic arm known for its high precision and 7-degree-of-freedom.
This setup enables us to simulate and evaluate complex tasks, and to define a common sense, which we showed in our paper how we learned and transfer it between tasks.
@@ Appendix B Experimental Details
Our velocity ground truth reward was calculated on the "end effector" part of the Sawyer robot shown in Fig. 9, and the end effector marked with an arrow. The end effector located at the end of the arm, there is an end effector, which can be equipped with different tools or grippers, depending on the task.
In simulation environments like MetaWorld, the end effector is often a gripper used for tasks like picking and placing objects.
The location of the end effector is given us by both Mujoco engine and Meta-worlds observation space. In each step we have an access to the end effector’s location of states $s$ and $s^{\prime}$, and we use those locations to calculate the velocity, as explained in section 5.
Figure 9: Sawyer Robot: The end effector, on which we calculated the velocity, is marked with an arrow.
We introduce two common sense behaviors, as shown in section 5. In the velocity case, we calculate the velocity of the Sawyer’s end effector. Our reward function is maximized when the end effector velocity get closer to $C_{v}$. In our experiments, we set $C_{v}$ to be 0.8. As for the action norm, our ground truth reward is maximized when the $l_{1}$ of the action space, gets closer to $C_{n}$.In our experiments, we set $C_{n}$ to be 3. It is important to mention that action space is the same for all meta worlds tasks.
Expert Demonstrations. In all our experiments, we used experts’ demonstration to train the discriminators. To create the expert demonstrations, we run SAC, and train in to maximize two reward functions. The first is the task-specific reward, and the second in the ground truth cs-rewards, as shown in Eq. 10 and 11. For SAC training, our hyperparameters (hp) are similar to the hp detailed in Meta-world benchmark. For hyperparameters values, see Table 5.
Code Base We use the garage toolkit, garage contributors [2019] which is a toolkit for developing and evaluating reinforcement learning algorithms, with a library of state-of-the-art implementations. We also utilized the Human-Compatible AI (HCAI) group’s project "Imitation Learning Baseline Implementations" Gleave et al. [2022] which provides clean implementations of imitation and reward learning algorithms.
   Description    Value
   Normal Hyperparameters
   Batch size    500
   Number of epochs    500
   Path length per roll-out    500
   Discount factor    0.99
   Algorithm-Specific Hyperparameters
   Policy hidden sizes    (256, 256)
   Activation function of hidden layers    ReLU
   Policy learning rate    $3\times 10^{-4}$
   Q-function learning rate    $3\times 10^{-4}$
   Policy minimum standard deviation    $e^{-20}$
   Policy maximum standard deviation    $e^{2}$
   Gradient steps per epoch    500
   Soft target interpolation parameter    $5\times 10^{-3}$
   Use automatic entropy tuning    True
Table 5: Hyperparameters used for Garage experiments with Single Task SAC
Experimental Details for CS-Reward Trained on a Single Task. The cs-reward is learned using Alg. 1, and we run in on the meta-worls envs: reach-wall, button-press-topdown-wall, drawer-close, push-back and coffee-button, separately. We train this experiment with $T=1$. another input is the expert demonstrations. We train the experts for this experiment as explained above in the current section. For each task, we sampled 60K trajectories. The Discriminator updates per iteration is 15, and the generator updates per iteration is 15. Our generator agent’s policy, and the new agent we trained from scratch with the learned cs-reward, are optimized with SAC algorithm. In Fig. 2, 4 and 5, Transferred graph and the SAC graph are averaged across five seeds.
Experimental Details for Learning CS-Reward with Multiple Targets. Experimental details are similar to experimental Details for CS-Reward Trained on a Single Task, except here we train Alg. 1 on multi-targets for each task. For each task, we train Alg. 1 on 5 different targets, means $T=5$. We learned a new cs-reward, and use it to train from scratch a new SAC agent on the same task, but with target. For each task, we train the SAC agent from scratch with the learned cs-reward on five different seeds. The results presented in Tables 1 and 3 are averaged across seeds.
Experimental Details for CS-Reward Learning CS-Reward with Multiple Tasks. In the Multiple Tasks tasks experiment, the setup is similar to the multi-target experiment, but here we train the Alg. 1, on multiple tasks. The differences are:
In this experiment, the expert demonstrations $\mathcal{D}_{t_{1}}...\mathcal{D}_{t_{T}}$, are set to 10K demonstrations for each task. Discriminator updates per iteration $D_{\text{updates}}$ is set to 10. Number of tasks $T$ is set to 5 targets * 6 tasks. The generator updates per iteration $G_{\text{updates}}$,is set to 20.
Experimental Details for Learning CS-Reward and Task-Reward with Multiple Tasks Here, the experimental details are similar to experimental details for "Learning CS-Reward with Multiple Targets", but in this experiment we train the Alg. 2, that contains two discriminators, and the number of discriminator updates per iteration is the same for both discriminators.
Compute Details: Algorithm 1 can be run on a single NVIDIA T4 GPU. For improved time performance, parallelization is recommended.